# Rodrigues Network for Learning Robot Actions

### v1

@@ Appendix A Rodrigues Network for animated characters
Beyond articulated robots, animated characters such as MANO [40] and SMPL [31] also exhibit articulated structures and employ forward kinematics, which can similarly serve as inductive biases for neural networks. However, a fundamental distinction exists between these models and articulated robots: each joint in MANO and SMPL is a free rotational joint with three degrees of freedom, as opposed to the single-axis, 1-DoF rotational joints typically found in robots. As a result, adapting our approach to animated characters requires a fundamental modification of the Rodrigues Operator to accommodate unconstrained 3-DoF rotations. We need to identify the basic operation in forward kinematics for animated characters, and develop a neural operator by turning that operation learnable.
@@
Appendix B Experiment settings
B.1 Forward kinematics fitting
@@
In this experiment, the models learn to fit the forward kinematics mapping of a robot. Therefore, we do not need to generate a fixed-size training set. However, we do create a fixed-size validation set for selecting the best checkpoint, and a separate test set for evaluation. We conduct this experiment on the LEAP hand, which is a fully-actuated robotic hand with four fingers, 16 rotational joints, and 17 links. Each data point includes the input: the root link position $T\in\mathbb{R}^{3}$, the root link orientation matrix $R\in\mathbb{R}^{3\times 3}$, and joint angles $\theta\in\mathbb{R}^{16}$. The output is the position vectors and orientation matrices of all 17 links. We sample $T$ uniformly in $[-0.05~{}\text{cm},0.05~{}\text{cm}]^{3}$, $R$ uniformly from $\mathrm{SO}(3)$, and $\theta$ uniformly within the joint limits. We then compute all link poses using forward kinematics. The validation and test sets each contain 10,000 input-output pairs, generated using different random seeds.
@@
B.2 Motion prediction in Cartesian space
@@
This experiment is conducted on a fixed-base 6-DoF UR5 robotic arm. To make the inverse kinematics solution unique for all reachable end-effector poses, we limit the six joint ranges to $[0,\pi/2]$, $[-\pi/2,0]$, $[0,\pi/2]$, $[0,\pi/4]$, $[0,\pi/4]$, and $[0,\pi/4]$. We sample two joint configurations $\boldsymbol{\theta}_{\text{start}},\boldsymbol{\theta}_{\text{end}}\in\mathbb{R}^{6}$ within these limits, which serve as the start and end poses. We then use forward kinematics to compute the corresponding 3D end-effector poses $\mathbf{P}_{\text{start}},\mathbf{P}_{\text{end}}\in\mathrm{SE}(3)$. Next, we interpolate $16-2$ intermediate 3D poses between the start and end to form a trajectory of 16 poses: $\mathbf{P}_{1},\cdots,\mathbf{P}_{16}$. Positions are interpolated linearly, and orientations are interpolated using spherical linear interpolation (slerp). Finally, we apply inverse kinematics to get the corresponding joint configurations $\boldsymbol{\theta}_{1},\cdots,\boldsymbol{\theta}_{16}$ for the 3D pose trajectory. The first 8 frames of joint configurations are used as input, and the models are asked to predict the last 8 frames. In Cartesian space, the motion pattern is clear: positions follow a straight line, and orientations change smoothly with slerp. However, this pattern is not obvious in joint space. The model needs to learn to reason about this motion pattern from the joint configurations. We generated four training sets with $10^{3}$, $10^{4}$, $10^{5}$, and $10^{6}$ input-output pairs. The validation and test sets each contain $10^{4}$ input-output pairs.
@@
We apply forward kinematics to the predicted joint configurations to obtain the predicted end-effector 3D poses. Then, we compare both the predicted joint configurations and the predicted end-effector poses with the ground truth. We report the following evaluation metrics: 1) $\text{Error}_{T}$ (mm): End-effector position error on the test set, averaged over the 8 predicted frames. 2) $\text{Error}_{R}$ (): End-effector orientation error on the test set, averaged over the 8 predicted frames. We compute the relative rotation between the predicted and ground truth orientations, convert it to axis-angle representation, and take the angle. 3) $\text{Error}_{\theta}$ (): Absolute joint configuration error on the test set, averaged over the 8 predicted frames and 6 joints. 4) MSE (×$\times 10^{-6}$): Mean squared error on the test set, multiplied by $10^{-6}$ for better readability. 5) Train MSE (×$\times 10^{-6}$): Mean squared error on the training set, also scaled by $10^{-6}$. The first four metrics measure the model’s generalization ability, while the last one reflects its fitting ability. For all metrics, lower values indicate better performance.
B.3 Robotic manipulation with imitation learning
@@
We evaluate on five representative manipulation tasks from the ManiSkill Benchmark [35]: PushCube, PickCube, StackCube, PegInsertionSide, and PlugCharger. For each task, we collect expert demonstration trajectories using the ManiSkill data collection API. These trajectories are generated via motion planning algorithms and are limited to a maximum of 200 simulation steps per episode. The number of demonstrations collected per task is listed in Table 7. During testing, the initial scene configurations are randomized using the same distribution employed during training data collection. For additional details, please refer to the ManiSkill benchmark [35].
@@
We follow the training setup in [14], including optimizer choice, learning rate scheduling, and exponential moving average (EMA) for network stabilization. Specifically, we use the AdamW optimizer with a learning rate of 0.0001, $\beta=(0.95,0.999)$, and a weight decay of $1\times 10^{-6}$. A cosine learning rate scheduler with 500 warm-up steps is applied. To enhance training speed and stability, we maintain an EMA of the model weights with a decay factor of 0.75. During training, we sample a batch of 1024 expert observation-action pairs from the demonstration dataset in each iteration and perform a gradient update using the diffusion loss. All training hyperparameters are summarized in Table 6. The number of training iterations and demonstration trajectories varies across tasks, depending on task complexity; see Table 7 for details. For further implementation details, please refer to [14].
Table 6: Training hyperparameters for imitation learning experiment (following [14]’s settings).
@@
Table 7: Demo trajectories and training iterations for each task in imitation learning experiment
@@
Following Hamer, we train our model on a large number of datasets that provide 2D or 3D hand annotations. Specifically, we use FreiHAND [59], HO3D [21], MTC [51], RHD [58], InterHand2.6M [33], H2O3D [21], DEX YCB [7], COCO WholeBody [25], Halpe [19] and MPII NZSL [44].
@@
We follow the training protocol described in Hamer [37]. We use the AdamW optimizer with learning rate 1e-5 and weight decay 1e-4, and train the model for 1,000,000 steps. The batch size is set to 64. Our model takes in a square image around the target hand, resized to 256 x 256, and output the 58 MANO [40] parameters of the human hand, specifically 48 pose parameters and the 10 shape parameters. We evaluate the model on the validation set for every 1,000 steps, and pick the model with the lowest validation loss for final evaluation. Similar to Hamer, we additionally estimate the camera parameters $\pi$. The camara $\pi$ corresponds to a translation $t\in\mathbb{R}^{3}$ that projects the 3D mesh and joints into the image. Hyperparameters are summarized in Table 8.
@@
Appendix C Method and implementation details
C.1 Forward kinematics fitting
We construct four baseline methods using existing neural network backbones for comparison: MLP, Graph Convolution Network (GCN), Transformer, and Body Transformer (BoT) [41].
@@ C.1 Forward kinematics fitting
C.2 Motion prediction in Cartesian space
@@
C.3 Robotic Manipulation with Imitation Learning
This experiment follows the Diffusion Policy (DP) framework proposed in [14], and evaluates different neural backbones for denoising action samples. The original paper [14] provides two backbone options, convolutional UNet and Transformer, which we directly adopt as baselines. In the DP framework, the policy observes a history of 2 prior time steps, predicts a sequence of 16 future actions, and executes the first 8 during deployment. Each action vector consists of 8 values: 7 for joint position control of the 7-DoF Franka arm and 1 for the 1-DoF Panda gripper.
@@ C.3 Robotic Manipulation with Imitation Learning
The UNet-DP baseline uses a convolutional UNet architecture that performs temporal convolution over the 16-step action sequence. We follow the default structure, with the hidden dimensions of the downsampling layers set to $[320,320,344]$. For further implementation details, please refer to [14].
@@
The Transformer-DP baseline applies causal self-attention to the 16 action-step tokens. We use an embedding dimension of 320, feed-forward hidden dimension of 1280, and a total of 10 Transformer blocks. Additional details are available in [14].
@@
C.4 3D hand reconstruction
Architecture-wise, we employ a Vision Transformer [17] as the image-processing backbone, followed by our Rodrigues Network (RodriNet) head to regress both hand and camera parameters. Since there are 17 links and 16 joints in the MANO parameter model, to make the information pass through the entire kinematic chain, our RodriNet head comprises 18 Rodrigues blocks, each of which sequentially applies a Rodrigues Layer, a Joint Layer, and a Self-attention Layer to update the link features, joint features. The Rodrigues Layer uses 4 link channels, and the Self-attention Layer uses 64 embed dimensions. We enable Global Tokens in this experiment to model the MANO’s shape parameters.
@@
Appendix D Additional results
@@ D.1 Ablation studies
Table 11: Ablation studies for motion prediction in Cartesian space with trainset size $=10^{5}$. We remove the Rodrigues Layer (R Layer), Joint Layer (J Layer), or Self-attention Layer (S Layer) respectively from the original Rodrigues Network, and evaluate the MSE on train/test sets.
@@
D.2 Tuning the baselines for motion prediction
Figure 6: Comparing our method to different baseline configurations in motion prediction with trainset size $=10^{5}$.
@@ D.2 Tuning the baselines for motion prediction
D.3 Hyperparameter sensitivity analysis for the Rodrigues Network
Table 12: Hyperparameter sensitivity analysis for the Rodrigues Network on motion prediction in Cartesian space with trainset size $=10^{5}$.
@@ D.3 Hyperparameter sensitivity analysis for the Rodrigues Network
Appendix E Accelerating the Multi-Channel Rodrigues Operator with CUDA

### v2

@@ Appendix A Rodrigues Network for animated characters
Beyond articulated robots, animated characters such as MANO (Romero et al., 2022) and SMPL (Loper et al., 2023) also exhibit articulated structures and employ forward kinematics, which can similarly serve as inductive biases for neural networks. However, a fundamental distinction exists between these models and articulated robots: each joint in MANO and SMPL is a free rotational joint with three degrees of freedom, as opposed to the single-axis, 1-DoF rotational joints typically found in robots. As a result, adapting our approach to animated characters requires a fundamental modification of the Rodrigues Operator to accommodate unconstrained 3-DoF rotations. We need to identify the basic operation in forward kinematics for animated characters, and develop a neural operator by turning that operation learnable.
@@
Appendix B Additional details on network architecture
This section describes how task observations are mapped to the Rodrigues Network’s inputs and how the network’s outputs are mapped to task actions.
Input embeddings
In all experiments except 3D hand reconstruction, task inputs are low-dimensional vectors. We use a linear input embedding layer to map the input observation feature to the Rodrigues Network’s input. Specifically, given an observation feature of dimension $d_{\text{obs}}$, we apply separate linear transformations to produce the initial link features, joint features, and, if applicable, a global token. These serve as inputs to the first Rodrigues Block of the Rodrigues Network. After processing through the stacked Rodrigues Blocks, we obtain output features for all links, joints, and the global token.
For 3D hand reconstruction, the inputs are RGB images. We first apply a ViT encoder to transform the image into visual tokens, then insert a cross-attention layer into each Rodrigues Block so that link features and the global token can attend to these tokens. As for the input to the first Rodrigues Block in this case, learnable embeddings are used.
Output heads
For joint-level outputs, each joint’s final output feature is concatenated with the output feature of its child link, then passed through a joint-specific linear transformation to produce the final output. For global outputs (those not associated with a specific joint), we apply a separate linear transformation to the final global token.
Appendix C Experiment settings
C.1 Forward kinematics fitting
@@
In this experiment, the models learn to fit the forward kinematics mapping of a robot. Therefore, we do not need to generate a fixed-size training set. However, we do create a fixed-size validation set for selecting the best checkpoint, and a separate test set for evaluation. We conduct this experiment on the LEAP hand, which is a fully-actuated robotic hand with four fingers, 16 rotational joints, and 17 links. Each data point includes the input: the root link position $T\in\mathbb{R}^{3}$, the root link orientation matrix $R\in\mathbb{R}^{3\times 3}$, and joint angles $\theta\in\mathbb{R}^{16}$. The output is the position vectors and orientation matrices of all 17 links. We sample $T$ uniformly in $0.05~\text{cm},0.05~\text{cm}]^{3}$, $R$ uniformly from $\mathrm{SO}(3)$, and $\theta$ uniformly within the joint limits. We then compute all link poses using forward kinematics. The validation and test sets each contain 10,000 input-output pairs, generated using different random seeds.
@@
C.2 Motion prediction in Cartesian space
@@
This experiment is conducted on a fixed-base 6-DoF UR5 robotic arm. To make the inverse kinematics solution unique for all reachable end-effector poses, we limit the six joint ranges to $0,\pi/2]$, $\pi/2,0]$, $0,\pi/2]$, $0,\pi/4]$, $0,\pi/4]$, and $0,\pi/4]$. We sample two joint configurations $\bm{\theta}_{\text{start}},\bm{\theta}_{\text{end}}\in\mathbb{R}^{6}$ within these limits, which serve as the start and end poses. We then use forward kinematics to compute the corresponding 3D end-effector poses $\mathbf{P}_{\text{start}},\mathbf{P}_{\text{end}}\in\mathrm{SE}(3)$. Next, we interpolate 16216-2 intermediate 3D poses between the start and end to form a trajectory of 16 poses: $\mathbf{P}_{1},\cdots,\mathbf{P}_{16}$. Positions are interpolated linearly, and orientations are interpolated using spherical linear interpolation (slerp). Finally, we apply inverse kinematics to get the corresponding joint configurations $\bm{\theta}_{1},\cdots,\bm{\theta}_{16}$ for the 3D pose trajectory. The first 8 frames of joint configurations are used as input, and the models are asked to predict the last 8 frames. In Cartesian space, the motion pattern is clear: positions follow a straight line, and orientations change smoothly with slerp. However, this pattern is not obvious in joint space. The model needs to learn to reason about this motion pattern from the joint configurations. We generated four training sets with $10^{3}$, $10^{4}$, $10^{5}$, and $10^{6}$ input-output pairs. The validation and test sets each contain $10^{4}$ input-output pairs.
@@
We apply forward kinematics to the predicted joint configurations to obtain the predicted end-effector 3D poses. Then, we compare both the predicted joint configurations and the predicted end-effector poses with the ground truth. We report the following evaluation metrics: 1) $\text{Error}_{T}$ (mm): End-effector position error on the test set, averaged over the 8 predicted frames. 2) $\text{Error}_{R}$ (): End-effector orientation error on the test set, averaged over the 8 predicted frames. We compute the relative rotation between the predicted and ground truth orientations, convert it to axis-angle representation, and take the angle. 3) $\text{Error}_{\theta}$ (): Absolute joint configuration error on the test set, averaged over the 8 predicted frames and 6 joints. 4) MSE ($\times 10^{-6}$): Mean squared error on the test set, multiplied by $10^{-6}$ for better readability. 5) Train MSE ($\times 10^{-6}$): Mean squared error on the training set, also scaled by $10^{-6}$. The first four metrics measure the model’s generalization ability, while the last one reflects its fitting ability. For all metrics, lower values indicate better performance.
C.3 Robotic manipulation with imitation learning
@@
We evaluate on five representative manipulation tasks from the ManiSkill Benchmark (Mu et al., 2021): PushCube, PickCube, StackCube, PegInsertionSide, and PlugCharger. For each task, we collect expert demonstration trajectories using the ManiSkill data collection API. These trajectories are generated via motion planning algorithms and are limited to a maximum of 200 simulation steps per episode. The number of demonstrations collected per task is listed in Table 7. During testing, the initial scene configurations are randomized using the same distribution employed during training data collection. For additional details, please refer to the ManiSkill benchmark (Mu et al., 2021).
@@
We follow the training setup in Chi et al. (2023), including optimizer choice, learning rate scheduling, and exponential moving average (EMA) for network stabilization. Specifically, we use the AdamW optimizer with a learning rate of 0.0001, $\beta=(0.95,0.999)$, and a weight decay of $1\times 10^{-6}$. A cosine learning rate scheduler with 500 warm-up steps is applied. To enhance training speed and stability, we maintain an EMA of the model weights with a decay factor of 0.75. During training, we sample a batch of 1024 expert observation-action pairs from the demonstration dataset in each iteration and perform a gradient update using the diffusion loss. All training hyperparameters are summarized in Table 6. The number of training iterations and demonstration trajectories varies across tasks, depending on task complexity; see Table 7 for details. For further implementation details, please refer to Chi et al. (2023).
Table 6: Training hyperparameters for imitation learning experiment (following Chi et al. (2023)’s settings).
@@
Table 7: Demo trajectories and training iterations for each task in imitation learning experiment.
@@
Following Hamer, we train our model on a large number of datasets that provide 2D or 3D hand annotations. Specifically, we use FreiHAND (Zimmermann et al., 2019), HO3D (Hampali et al., 2020), MTC (Xiang et al., 2019), RHD (Zimmermann and Brox, 2017), InterHand2.6M (Moon et al., 2020), H2O3D (Hampali et al., 2020), DEX YCB (Chao et al., 2021), COCO WholeBody (Jin et al., 2020), Halpe (Fang et al., 2022) and MPII NZSL (Simon et al., 2017).
@@
We follow the training protocol described in HaMeR (Pavlakos et al., 2024). We use the AdamW optimizer with learning rate 1e-5 and weight decay 1e-4, and train the model for 1,000,000 steps. The batch size is set to 64. Our model takes in a square image around the target hand, resized to 256 x 256, and output the 58 MANO (Romero et al., 2022) parameters of the human hand, specifically 48 pose parameters and the 10 shape parameters. We evaluate the model on the validation set for every 1,000 steps, and pick the model with the lowest validation loss for final evaluation. Similar to Hamer, we additionally estimate the camera parameters $\pi$. The camara $\pi$ corresponds to a translation $t\in\mathbb{R}^{3}$ that projects the 3D mesh and joints into the image. Hyperparameters are summarized in Table 8.
@@
Appendix D Method and implementation details
D.1 Forward kinematics fitting
We construct four baseline methods using existing neural network backbones for comparison: MLP, Graph Convolution Network (GCN), Transformer, and Body Transformer (BoT) (Sferrazza et al., 2024).
@@ D.1 Forward kinematics fitting
D.2 Motion prediction in Cartesian space
@@
D.3 Robotic Manipulation with Imitation Learning
This experiment follows the Diffusion Policy (DP) framework proposed in Chi et al. (2023), and evaluates different neural backbones for denoising action samples. The original paper (Chi et al., 2023) provides two backbone options, convolutional UNet and Transformer, which we directly adopt as baselines. In the DP framework, the policy observes a history of 2 prior time steps, predicts a sequence of 16 future actions, and executes the first 8 during deployment. Each action vector consists of 8 values: 7 for joint position control of the 7-DoF Franka arm and 1 for the 1-DoF Panda gripper.
@@ D.3 Robotic Manipulation with Imitation Learning
The UNet-DP baseline uses a convolutional UNet architecture that performs temporal convolution over the 16-step action sequence. We follow the default structure, with the hidden dimensions of the downsampling layers set to $[320,320,344]$. For further implementation details, please refer to Chi et al. (2023).
@@
The Transformer-DP baseline applies causal self-attention to the 16 action-step tokens. We use an embedding dimension of 320, feed-forward hidden dimension of 1280, and a total of 10 Transformer blocks. Additional details are available in Chi et al. (2023).
@@
D.4 3D hand reconstruction
Architecture-wise, we employ a Vision Transformer (Dosovitskiy et al., 2020) as the image-processing backbone, followed by our Rodrigues Network (RodriNet) head to regress both hand and camera parameters. Since there are 17 links and 16 joints in the MANO parameter model, to make the information pass through the entire kinematic chain, our RodriNet head comprises 18 Rodrigues blocks, each of which sequentially applies a Rodrigues Layer, a Joint Layer, and a Self-attention Layer to update the link features, joint features. The Rodrigues Layer uses 4 link channels, and the Self-attention Layer uses 64 embed dimensions. We enable Global Tokens in this experiment to model the MANO’s shape parameters.
@@
Appendix E Additional results
@@ E.1 Ablation studies
Table 11: Ablation studies for motion prediction in Cartesian space with trainset size $10^{5}$. We remove the Rodrigues Layer (R Layer), Joint Layer (J Layer), or Self-attention Layer (S Layer) respectively from the original Rodrigues Network, and evaluate the MSE on train/test sets.
@@
E.2 Tuning the baselines for motion prediction
Figure 6: Comparing our method to different baseline configurations in motion prediction with trainset size $10^{5}$.
@@ E.2 Tuning the baselines for motion prediction
E.3 Hyperparameter sensitivity analysis for the Rodrigues Network
Table 12: Hyperparameter sensitivity analysis for the Rodrigues Network on motion prediction in Cartesian space with trainset size $10^{5}$.
@@ E.3 Hyperparameter sensitivity analysis for the Rodrigues Network
Appendix F Accelerating the Multi-Channel Rodrigues Operator with CUDA